OpenStudy (anonymous):

Which of the following points is a solution to the following system of equations? x + y = 1 3x – y = 3 A. (3, 0) B. (5, –1) C. (1, 4) D. (1, 0)

OpenStudy (anonymous):

@texaschic101 @amistre64

OpenStudy (texaschic101):

look at your first equation... x + y = 1 now look at your answer choices...which answer choice fits into this equation and makes it true ?

OpenStudy (texaschic101):

does 3 + 0 = 1 ? Does 5 + (-1) = 1 ?? Does 1 + 4 = 1 ? Does 1 + 0 = 1 ? Come on...this is not that hard

OpenStudy (texaschic101):

once you find the answers choice that makes your first equation true, then sub that into the 2nd equation and see if it satisfies that one as well. If so, then you have just found your answer